Calculate the mass of the product of 6.40 g of magnesium with 1.32 g of oxygen . 



1
Expert's answer
2021-02-01T03:47:07-0500

The balanced reaction equation is:

2Mg + O2 "\rightarrow" 2MgO.


As one can see from the equation, 2 moles of magnesium react with 1 mol of oxygen and 2 moles of magnesium oxide are formed:

"\frac{n(Mg)}{2} = n(O_2)=\frac{n(MgO)}{2}" .


The number of the moles is the mass "m" divided by the molar mass "M" :

"n = \frac{m}{M}" .


The molar masses of Mg, O2 and MgO are 24.31, 32.00 and 40.30 g/mol, respectively. Therefore, the number of the moles of Mg are O2 are:

"n(Mg) = \frac{6.40}{24.31} = 0.2633" mol.

"n(O_2) = \frac{1.32}{32.00} = 0.04125" mol.


The number of the moles of Mg divided by two 0.2633/2 = 0.1316 mol is higher than the number of the moles of O2. Thus, Mg is in excess. Therefore, the quantity of the product, MgO, must be calculated from the number of the moles of O2:

"n(MgO) = 2n(O_2) = 2\u00b70.04125 = 0.0825" mol.


Finally, the mass of magnesium oxide formed is:

"m(MgO) =n\u00b7M = 0.0825\u00b740.30 = 3.32" g.


Answer: the mass of the product of 6.40 g of magnesium with 1.32 g of oxygen is 3.32 g.